# Cohomology of Lie algebroids over Algebraic spaces

### v6

@@ 5.2. Hochschild hypercohomology of universal enveloping algebroid
(Noncommutative analogue of HKR Theorem) The free Lie algebroid $\mathcal{P}_{X}$ can be identified with sheaf of regular noncommutative vector fields on $X$ and we have seen already that the sections of its universal enveloping algebra $\mathbb{D}_{X}$ are noncommutative differential operators on $X$ (see Remark 2.12). Thus, in this case from the HKR theorem we get a canonical isomorphism
@@
For a cocomplete graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra $\mathcal{A}$ (of finite or infinite type), the isomorphism $\mathcal{A}\cong\mathscr{U}(\mathcal{O}_{X},\mathscr{P}(\mathcal{A}))$ holds [MS22] and thus we get the graded vector space isomorphism
@@ 5.3. Hochschild hypercohomology of jet algebroid
Here, we present a dual version of the HKR theorem by considering Hochschild cohomology of a jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ for a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ (see [MS22]).
@@
There is a canonical left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module structure on $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ constructed as follows (see [CVdB10, KP11, CRV12] for the local descriptions):
@@
(Generalized dual HKR theorem) Let $\mathcal{L}$ be a locally free Lie algebroid over $(X,\mathcal{O}_{X})$ which is of finite rank. Then the anti-symmetrization map (dual HKR morphism)
@@
For a locally free Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ of finite rank (say $r$, for some $r\in\mathbb{N}$), the dual of $\mathcal{L}$, denoted by $\mathcal{L^{*}}$, is a locally free $\mathcal{O}_{X}$-module of the same rank. Thus, $\mathcal{L^{*}}(U_{x})$ has a basis $\{w_{1},\dots,w_{r}\}$ (say) for each special open set $U_{x}$ around $x\in X$. Since $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ is a commutative $\mathcal{O}_{X}$-algebra, thus $\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\cong\mathcal{O}_{X}(U_{x})[[w_{1},\dots,w_{r}]]$ as $\mathcal{O}_{X}(U_{x})$-algebra. Hence, by applying sheafification functor we get the isomorphism of $\mathcal{O}_{X}$-algebras as $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\cong\widehat{\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L^{*}}}$ (this sheaf of symmetric algebras is formally completed with respect to the degree) [CRvdB10, BP15]. Now, consider the dual Koszul-Rinehart resolution of $\mathcal{O}_{X}$ in the category of left $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$-comodules (by using local counterpart from [KP11]) is
@@
Applying cotensor product by $\mathcal{O}_{X}$ (with canonical $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$-comodule structure, which induces from the $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module structure on $\mathcal{O}_{X}$) to the dual Koszul-Rinehart complex $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$, we get the canonical isomorphism (sheafifying local descriptions from [KP11]) of cochain complexes of sheaves of graded vector spaces as
@@
Since $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module of finite rank, there exists open sets $U_{x}$ around each point $x\in X$, where $\mathcal{L}|_{U_{x}}$ is free $\mathcal{O}_{X}|_{U_{x}}$-module of finite rank. Thus, for each $U_{x}$ we can use results from the proof of the dual HKR Theorem [KP11] for the $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$. Here, we get
@@
In [CRvdB10, CRV12], a dual version of the (twisted) HKR theorem for Lie algebroids [CVdB10] (with rich algebraic structures) is considered in-order to study a notion of precalculus upto homotopy in the Lie algebroids context.
@@
In the special case when $\mathcal{L}=\mathcal{T}_{X}$ and $X$ is non-singular (smooth manifold, complex manifold, smooth algebraic variety or smooth scheme over the field $\mathbb{C})$, we get isomorphism of graded vector spaces
@@
by applying de Rham theorems in different settings (smooth, analytic, algebraic) [Dri06, Ste15], where $\mathbb{K=R}$ or $\mathbb{C}$ accordingly and $X^{an}$ is the analytification of the algebraic variety (scheme) $X$.
@@
Applying the Theorem 3.5 for $\mathcal{L}=\mathcal{T}_{X}$ over some non-singular space $X$ and using the Corollary 5.14 we get the isomorphism
@@
where $\mathcal{J}_{X}:=\mathscr{J}(\mathcal{O}_{X},\mathcal{T}_{X})$ is the usual sheaf of jets on $X$ [CRvdB10, Pym13].
Acknowledgment. I wholeheartedly express my deepest gratitude to my PhD thesis supervisor Prof. Ashis Mandal for his immense support and guidance.

### v7

@@ 5.2. Hochschild hypercohomology of universal enveloping algebroid
((((Noncommutative analogue of HKR Theorem)))) The free Lie algebroid $\mathcal{P}_{X}$ can be identified with sheaf of regular noncommutative vector fields on $X$ and we have seen already that the sections of its universal enveloping algebra $\mathbb{D}_{X}$ are noncommutative differential operators on $X$ ((((see Remark 2.12)))). Thus, in this case from the HKR theorem we get a canonical isomorphism
@@
For a cocomplete graded free $\mathcal{O}_{X}/{\mathbb{K}_{X}}$-bialgebra $\mathcal{A}$ ((((of finite or infinite type)))), the isomorphism $\mathcal{A}\cong\mathscr{U}(\mathcal{O}_{X},\mathscr{P}(\mathcal{A}))$ holds [29] and thus we get the graded vector space isomorphism
@@ 5.3. Hochschild hypercohomology of jet algebroid
Here, we present a dual version of the HKR theorem by considering Hochschild cohomology of a jet algebroid $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ for a Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ (see the part (2.13) mentioned in Section 2).
@@
There is a canonical left $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module structure on $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ constructed as follows ((((see [7, 19, 6] for the local descriptions)))):
@@
((((Generalized dual HKR theorem)))) Let $\mathcal{L}$ be a locally free Lie algebroid over $(X,\mathcal{O}_{X})$ which is of finite rank. Then the anti-symmetrization map ((((dual HKR morphism))))
@@
For a locally free Lie algebroid $\mathcal{L}$ over $(X,\mathcal{O}_{X})$ of finite rank (say $r$, for some $r\in\mathbb{N}$), the dual of $\mathcal{L}$, denoted by $\mathcal{L^{*}}$, is a locally free $\mathcal{O}_{X}$-module of the same rank. Thus, $\mathcal{L^{*}}(U_{x})$ has a basis $\{w_{1},\dots,w_{r}\}$ (say) for each special open set $U_{x}$ around $x\in X$. Since $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ is a commutative $\mathcal{O}_{X}$-algebra, thus $\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\cong\mathcal{O}_{X}(U_{x})[[w_{1},\dots,w_{r}]]$ as $\mathcal{O}_{X}(U_{x})$-algebra. Hence, by applying sheafification functor we get the isomorphism of $\mathcal{O}_{X}$-algebras as $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})\cong\widehat{\mathcal{S}_{\mathcal{O}_{X}}\mathcal{L^{*}}}$ (this sheaf of symmetric algebras is formally completed with respect to the degree) [5, 2]. Now, consider the dual Koszul-Rinehart resolution of $\mathcal{O}_{X}$ in the category of left $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$-comodules (by using local counterpart from [19]) is
@@
Applying cotensor product by $\mathcal{O}_{X}$ (with canonical $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$-comodule structure, which induces from the $\mathscr{U}(\mathcal{O}_{X},\mathcal{L})$-module structure on $\mathcal{O}_{X}$) to the dual Koszul-Rinehart complex $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$, we get the canonical isomorphism (sheafifying local descriptions from [19]) of cochain complexes of sheaves of graded vector spaces as
@@
Since $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module of finite rank, there exists open sets $U_{x}$ around each point $x\in X$, where $\mathcal{L}|_{U_{x}}$ is free $\mathcal{O}_{X}|_{U_{x}}$-module of finite rank. Thus, for each $U_{x}$ we can use results from the proof of the dual HKR Theorem [19] for the $(\mathbb{K},\mathcal{O}_{X}(U_{x}))$-Lie-Rinehart algebra $\mathcal{L}(U_{x})$. Here, we get
@@
In [5, 6], a dual version of the (twisted) HKR theorem for Lie algebroids [7] (with rich algebraic structures) is considered in-order to study a notion of precalculus upto homotopy in the Lie algebroids context.
@@
In particular, when $\mathcal{L}=\mathcal{T}_{X}$ over a non-singular $a$-space (smooth manifold, complex manifold, smooth algebraic variety or smooth scheme over the field $\mathbb{C})$ $X$, we get isomorphism of graded vector spaces
@@
by applying de Rham theorems in different settings ((((smooth, analytic, algebraic)))) [9, 36], where $\mathbb{K=R}$ or $\mathbb{C}$ accordingly and $X^{an}$ is the analytification of the algebraic variety (scheme) $X$.
@@
Applying the Theorem 3.6 for $\mathcal{L}=\mathcal{T}_{X}$ over some non-singular $a$-space $X$ and using the Corollary 5.14 we get the isomorphism
@@
where $\mathcal{J}_{X}:=\mathscr{J}(\mathcal{O}_{X},\mathcal{T}_{X})$ is the usual sheaf of jets on $X$ [5, 32].
5.4. Remarks on Lie algebroids over Schemes.
All the previous discussions are applicable to Lie algebroids over noetherian separated schemes (or schemes of finite types) over a field of characteristic zero.